# Playmate2: Training-Free Multi-Character Audio-Driven Animation via Diffusion Transformer with Reward Feedback

### v1

@@ Abstract
Figure 1: We present a novel DiT-based framework for generating high-quality, audio-driven human videos that effectively tackles key challenges related to temporal coherence in long sequences and multi-character animations. To the best of our knowledge, this is the first training-free approach capable of enabling audio-driven animation for three or more characters without requiring additional data or model modifications.
@@ 1 Introduction
Benefiting from large-scale pre-training and advanced architectural designs, diffusion models (Rombach et al. 2022; Podell et al. 2023; Esser et al. 2024; Liu et al. 2024; Lin et al. 2024a; Yang et al. 2024; Kong et al. 2024; Wan et al. 2025) have achieved significant advances in image and video synthesis, outperforming traditional generative adversarial networks (GANs) (Goodfellow et al. 2020) in both visual quality and temporal coherence. These advancements have greatly improved the generation of audio-driven human videos (Xue et al. 2024), making this capability a cornerstone of digital human research. Audio-driven human animation has broad applications in digital entertainment, film and gaming production, virtual reality, and digital storytelling.
@@
Audio-driven human animation (Jiang et al. 2024) synthesizes realistic character videos with synchronized lip movements and natural body gestures from speech and auxiliary inputs. Recent advances in diffusion models have spurred their use in this domain, leading to two main categories: portrait animation and human animation. The first focuses on synthesizing facial expressions solely from audio signals, with little attention given to background dynamics (Tian et al. 2024; Xu et al. 2024a, b; Cui et al. 2024; Zheng et al. 2024; Ji et al. 2024; Ma et al. 2023b; Chen et al. 2024; Cui et al. 2025b). Such a restricted approach frequently compromises the realism of generated videos in complex scenes, leading to results that do not satisfy the demands of high-quality applications. The second employs video diffusion models to overcome the aforementioned spatial constraints, thereby achieving full-body animation generation (Lin et al. 2025; Fei et al. 2025; Wang et al. 2025a; Chen et al. 2025; Kong et al. 2025; Cui et al. 2025a). Despite progress, several challenges persist: 1) Existing methods often struggle to maintain accurate lip-sync while generating natural body movements; 2) In long video synthesis, current solutions often result in jittery motions and abrupt transitions, failing to preserve temporal coherence; 3) Most existing techniques are unable to animate scenes involving multiple characters using audio input; although some works achieve multi-character animation by constructing multi-speaker datasets and introducing significant modifications to the model architecture, such strategies are often resource-intensive and not scalable.
@@
To address these challenges, leveraging the large-scale video diffusion model Wan2.1 (Wan et al. 2025), we propose a diffusion transformer (DiT)-based (Peebles and Xie 2023) framework for audio-driven facial and human video generation, aiming to enhance video quality and enable cost-effective multi-character animation. First, we adopt a LoRA-based (Hu et al. 2022) training strategy to preserve the capabilities of the foundation model while enabling long video generation. Next, we explore a training strategy that combines partial parameter updates with reward feedback, producing videos with accurate lip synchronization and natural body motions. Finally, inspired by Classifier-Free Guidance (CFG) (Ho and Salimans 2022), we introduce a training-free approach, Mask Classifier-Free Guidance (Mask-CFG), to tackle the challenge of multi-character animation. This method does not require constructing specialized datasets or modifying the model architecture; instead, it achieves multi-character control through simple adjustments during inference, making it both efficient and cost-effective.
@@ 2.1 Audio-Driven Portrait Animation
Prior work on audio-driven portrait animation has largely focused on lip-sync accuracy (Prajwal et al. 2020; Zhang et al. 2023b, a; Ma et al. 2023a; Guo et al. 2021; Wang et al. 2024; Ye et al. 2023). Traditional approaches based on GANs, neural radiance fields (NeRF) (Mildenhall et al. 2021), and 3D Gaussian Splatting (Kerbl et al. 2023) have achieved strong results, yet often fail to model the subtle relationship between prosody and facial dynamics, leading to limited expressiveness and reduced visual realism. Recently, diffusion-based methods have enabled end-to-end talking video generation. EMO (Tian et al. 2024) improves inter-frame consistency for stable, natural synthesis. Hallo (Xu et al. 2024a) jointly addresses lip synchronization, expression, and pose. Sonic (Ji et al. 2024) emphasizes global perceptual coherence for diverse motions, while DICE-Talk (Tan et al. 2025) and Playmate (Ma et al. 2025) introduce emotional control for expressive portraits. Despite producing realistic outputs, these methods are primarily limited to facial animation and do not support full-body motion synthesis.
@@ 2.2 Audio-Driven Human Animation
To enable audio-driven human animation, recent methods leverage large-scale video diffusion models. CyberHost (Lin et al. 2024b) proposes a one-stage framework with novel attention and human-prior-guided training for upper-body synthesis. Approaches like OmniHuman-1 (Lin et al. 2025), FantasyTalking (Wang et al. 2025a), SkyReels-Audio (Fei et al. 2025), and OmniAvatar (Gan et al. 2025) build on models such as Seaweed (Seawead et al. 2025), HunyuanVideo (Kong et al. 2024), and Wan2.1 for holistic motion generation. In multi-character scenarios, HunyuanVideo-Avatar (Chen et al. 2025) uses latent-space masking for localized, character-specific control, while MultiTalk (Kong et al. 2025) introduces Label Rotary Position Embedding with a multi-person dataset to resolve audio-person binding. Inspired by these advances, we base our approach on a large-scale video diffusion transformer for audio-driven human animation.
@@ 2.3 Direct Preference Optimization
Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al. 2022) is widely used to align large language models with human preferences. This paradigm has been extended to image and video generation via reward models or preference data (Zhang et al. 2024; Wallace et al. 2024; Liu et al. 2025). Recently, Direct Preference Optimization (DPO) (Rafailov et al. 2023) has gained traction in audio-driven animation. Hallo4 (Cui et al. 2025a) proposes a DPO framework for human-centric animation, leveraging a curated human preference dataset to align generated outputs with perceptual metrics related to motion-video alignment and facial expression naturalness. EchoMimicV3 (Meng et al. 2025) further adopts an alternating Supervised Fine-Tuning (SFT) and DPO training paradigm, enabling high-quality video generation with a 1.3B-parameter model. Building on these advances, we present a more efficient framework that integrates DPO to simultaneously enhance lip synchronization accuracy and facial expression naturalness in audio-driven video generation.
@@ 3.1 LoRA-based Long Video Generation
The framework of our method is illustrated in Figure 2, where Wan2.1 serves as the foundational model. Specifically, we employ the causal 3D Variational Autoencoder (VAE) (Kingma, Welling et al. 2013) from Wan2.1 to compress both the reference image and the ground-truth video from pixel space to the latent space. Additionally, we use UMT5 (Chung et al. 2023) for text encoding and CLIP (Radford et al. 2021) for image encoding. For audio input, we utilize Wav2Vec (Baevski et al. 2020) to extract audio tokens containing rich multi-scale acoustic features, which are then injected into the DiT through cross-attention mechanisms.
HunyuanVideo-Avatar (Chen et al. 2025) uses the Time-Aware Position Shift Fusion method from Sonic (Ji et al. 2024) to enable long video generation. OmniAvatar (Gan et al. 2025) reuses the final latent of the current segment as the initial latent for the next, and applies reference image embedding to preserve identity and maintain frame overlap for temporal consistency. Our experiments show that these two methods fail to achieve satisfactory performance in long video generation. This issue stems from the special architecture of video diffusion models, such as Wan2.1, which are designed to support joint training on both video and image data. In particular, given an input video $V\in\mathbb{R}^{(1+T)\times H\times W\times 3}$, where the frames of $V$ follow the $1+T$ input format, Wan2.1 divides the video into $1+T/4$ chunks. Then, Wan-VAE compresses the spatio-temporal dimensions of these chunks to $[1+T/4,H/8,W/8]$, while the first frame is only spatially compressed to better handle image data. This independent processing of the first frame tends to cause forgetting and drifting issues.
@@ 3.2 Partial-update Training and DPO
After completing the first LoRA-based training stage, we obtain a diffusion transformer capable of seamless long video generation. Next, we introduce the Audio Cross-Attention module and adopt the Flow Matching (Lipman et al. 2022) approach used in Wan2.1 to update its parameters. Specifically, we aggregate every four consecutive audio frames into a single representation to ensure temporal alignment between the audio features and the compressed video latent representation. The Audio Cross-Attention mechanism is defined as:
@@
To further improve lip-sync accuracy and align the model with human preferences, we introduce DPO for optimization after completing the aforementioned stages. Hallo4 (Cui et al. 2025a) presents the first audio-driven portrait DPO dataset that captures human preferences in lip-sync and facial naturalness via annotator rankings. Unlike Hallo4, which relies on human annotators to construct the dataset, we introduce DPO in a more efficient and cost-effective manner.
Direct Preference Optimization formulates the alignment of models with human preferences as a policy optimization task, based on pairwise preference data $\mathcal{D}=\{(x,y^{w},y^{l})\}$, where $y^{w}$ is preferred over $y^{l}$. As shown in Figure 2, for each training sample, we randomly select five segments and employ LatentSync (Li et al. 2024) to compute the Sync-C score for each; the highest-scoring segment is selected as $y^{w}$, and the lowest as $y^{l}$. Finally, we use the Flow-DPO loss proposed by VideoReward (Liu et al. 2025) to train the model:
@@ 3.3 Mask-CFG for Multi-Character Audio-driven Animation
After the training stage described above, we obtain a diffusion transformer that achieves accurate lip-sync and strong alignment with human preferences. We now introduce a training-free method to enable multi-character audio-driven video generation. Methods such as MultiTalk (Kong et al. 2025) and HunyuanVideo-Avatar achieve this capability by constructing multi-speaker datasets and modifying the cross-attention mechanism. In contrast, we enable multi-character animation by improving the classifier-free guidance (CFG) mechanism during inference—without any training or model modification—resulting in a simple, efficient, and framework-agnostic approach. Specifically, we propose Mask-CFG, which leverages spatial masks to route audio conditions to specific characters. Given an audio condition set $A=\{a_{1},a_{2},\dots,a_{n}\}$, the corresponding binary mask set is defined as $M=\{m_{1},m_{2},\dots,m_{n}\}$. Here, $a_{1}$ is considered to be silent audio, and $m_{1}$ serves as the background mask. Each $m_{i}\in\{0,1\}^{H\times W}$ denotes a binary segmentation of the input image, and the masks are exhaustive and mutually exclusive, satisfying $\bigvee_{i=1}^{n}m_{i}=\mathbf{1}$, meaning their union covers the entire image region. Under classifier-free guidance, the conditional distribution $p(a_{i}\mid x_{t})$ leads to the following formulation:
@@ 4.1 Experimental Setups
We collect our training data from public datasets (including AVSpeech (Ephrat et al. 2018) and OpenHumanVid (Li et al. 2025)) and sources we collect ourselves. To ensure high data quality, we employ tools such as Koala-36M (Wang et al. 2025b) to filter out videos with low brightness or poor aesthetic quality. Through this standardized selection process, we obtain over $300,000$ training samples, with a total duration exceeding $800$ hours. To demonstrate the effectiveness of our multi-character audio-driven approach in a training-free manner, all training samples are single-person talking videos. For evaluation, we use two public datasets: CelebV-HQ (Zhu et al. 2022), which features diverse scenes, and HDTF (Zhang et al. 2021), which provides high-resolution videos and a larger number of subjects, to assess the animation capabilities of our method.
@@
The training weights of our first LoRA stage are initialized from the pretrained Wan2.1-I2V-14B-720P model, and the training process is conducted using $16$ NVIDIA A100 GPUs for a total of $5000$ steps. Subsequently, we use $32$ NVIDIA A100 GPUs and conduct additional training for $100,000$ steps to obtain the first $v_{\text{ref}}$. Next, we introduce DPO-based refinement training for another $100,000$ steps, during which $v_{\text{ref}}$ is updated every $10,000$ steps. The model operates at a resolution of $720\times 1280$. We employ AdamW as the optimizer and set the learning rate to $1\times 10^{-5}$. After completing the above training pipeline, we introduce Mask CFG during the inference stage to enable multi-person audio-driven animation, with $\lambda$ set to $5.0$.

### v2

@@ 1 Introduction
Benefiting from large-scale pre-training and advanced architectural designs, diffusion models (rombach2022high; esser2024scaling; liu2024sora; yang2024cogvideox; kong2024hunyuanvideo; wan2025wan) have achieved significant advances in image and video synthesis, outperforming traditional generative adversarial networks (GANs) (goodfellow2020generative) in both visual quality and temporal coherence. These advancements have greatly improved the generation of audio-driven human videos (xue2024human), making this capability a cornerstone of digital human research. Audio-driven human animation has broad applications in digital entertainment, film and gaming production, virtual reality, and digital storytelling.
@@
Audio-driven human animation (jiang2024audio) synthesizes realistic character videos with synchronized lip movements and natural body gestures from speech and auxiliary inputs. Recent advances in diffusion models have spurred their use in this domain, leading to two main categories: portrait animation and human animation. The first focuses on synthesizing facial expressions solely from audio signals, with little attention given to background dynamics (tian2024emo; xu2024hallo; xu2024vasa; ji2024sonic; ma2023dreamtalk; chen2024echomimic; cui2025hallo3). Such a restricted approach frequently compromises the realism of generated videos in complex scenes, leading to results that do not satisfy the demands of high-quality applications. The second employs video diffusion models to overcome the aforementioned spatial constraints, thereby achieving full-body animation generation (lin2025omnihuman; fei2025skyreels; wang2025fantasytalking; chen2025hunyuanvideo; kong2025let; cui2025hallo4). Despite progress, several challenges persist: 1) Existing methods often struggle to maintain accurate lip-sync while generating natural body movements; 2) In long video synthesis, current solutions often result in jittery motions and abrupt transitions, failing to preserve temporal coherence; 3) Most existing techniques are unable to animate scenes involving multiple characters using audio input; although some works achieve multi-character animation by constructing multi-speaker datasets and introducing significant modifications to the model architecture, such strategies are often resource-intensive and not scalable.
@@
To address these challenges, leveraging the large-scale video diffusion model Wan2.1 (wan2025wan), we propose a diffusion transformer (DiT)-based (peebles2023scalable) framework for audio-driven facial and human video generation, aiming to enhance video quality and enable cost-effective multi-character animation. First, we adopt a LoRA-based (hu2022lora) training strategy to preserve the capabilities of the foundation model while enabling long video generation. Next, we explore a training strategy that combines partial parameter updates with reward feedback, producing videos with accurate lip synchronization and natural body motions. Finally, inspired by Classifier-Free Guidance (CFG) (ho2022classifier), we introduce a training-free approach, Mask Classifier-Free Guidance (Mask-CFG), to tackle the challenge of multi-character animation. This method does not require constructing specialized datasets or modifying the model architecture; instead, it achieves multi-character control through simple adjustments during inference, making it both efficient and cost-effective.
@@ 2.1 Audio-Driven Portrait Animation
Prior work on audio-driven portrait animation has largely focused on lip-sync accuracy (prajwal2020lip; zhang2023dinet; zhang2023sadtalker; guo2021ad; wang2024expression). Traditional approaches based on GANs, neural radiance fields (NeRF) (mildenhall2021nerf), and 3D Gaussian Splatting (kerbl20233d) have achieved strong results, yet often fail to model the subtle relationship between prosody and facial dynamics, leading to limited expressiveness and reduced visual realism. Recently, diffusion-based methods have enabled end-to-end talking video generation. EMO (tian2024emo) improves inter-frame consistency for stable, natural synthesis. Hallo (xu2024hallo) jointly addresses lip synchronization, expression, and pose. Sonic (ji2024sonic) emphasizes global perceptual coherence for diverse motions, while DICE-Talk (tan2025disentangle) and Playmate (maplaymate) introduce emotional control for expressive portraits. Despite producing realistic outputs, these methods are primarily limited to facial animation and do not support full-body motion synthesis.
@@ 2.2 Audio-Driven Human Animation
To enable audio-driven human animation, recent methods leverage large-scale video diffusion models. CyberHost (lin2024cyberhost) proposes a one-stage framework with novel attention and human-prior-guided training for upper-body synthesis. Approaches like OmniHuman-1 (lin2025omnihuman), FantasyTalking (wang2025fantasytalking), SkyReels-Audio (fei2025skyreels), and OmniAvatar (gan2025omniavatar) build on models such as Seaweed (seawead2025seaweed), HunyuanVideo (kong2024hunyuanvideo), and Wan2.1 for holistic motion generation. In multi-character scenarios, HunyuanVideo-Avatar (chen2025hunyuanvideo) uses latent-space masking for localized, character-specific control, while MultiTalk (kong2025let) introduces Label Rotary Position Embedding with a multi-person dataset to resolve audio-person binding. Inspired by these advances, we base our approach on a large-scale video diffusion transformer for audio-driven human animation.
@@ 2.3 Direct Preference Optimization
Reinforcement Learning from Human Feedback (RLHF) (ouyang2022training) is widely used to align large language models with human preferences. This paradigm has been extended to image and video generation via reward models or preference data (zhang2024direct; wallace2024diffusion; liu2025improving). Recently, Direct Preference Optimization (DPO) (rafailov2023direct) has gained traction in audio-driven animation. Hallo4 (cui2025hallo4) proposes a DPO framework for human-centric animation, leveraging a curated human preference dataset to align generated outputs with perceptual metrics related to motion-video alignment and facial expression naturalness. EchoMimicV3 (meng2025echomimicv3) further adopts an alternating Supervised Fine-Tuning (SFT) and DPO training paradigm, enabling high-quality video generation with a 1.3B-parameter model. Building on these advances, we present a more efficient framework that integrates DPO to simultaneously enhance lip synchronization accuracy and facial expression naturalness in audio-driven video generation.
@@ 3.1 LoRA-based Long Video Generation
The framework of our method is illustrated in Figure 2, where Wan2.1 serves as the foundational model. Specifically, we employ the causal 3D Variational Autoencoder (VAE) (kingma2013auto) from Wan2.1 to compress both the reference image and the ground-truth video from pixel space to the latent space. Additionally, we use UMT5 (chung2023unimax) for text encoding and CLIP (radford2021learning) for image encoding. For audio input, we utilize Wav2Vec (baevski2020wav2vec) to extract audio tokens containing rich multi-scale acoustic features, which are then injected into the DiT through cross-attention mechanisms.
HunyuanVideo-Avatar (chen2025hunyuanvideo) uses the Time-Aware Position Shift Fusion method from Sonic (ji2024sonic) to enable long video generation. OmniAvatar (gan2025omniavatar) reuses the final latent of the current segment as the initial latent for the next, and applies reference image embedding to preserve identity and maintain frame overlap for temporal consistency. Our experiments show that these two methods fail to achieve satisfactory performance in long video generation. This issue stems from the special architecture of video diffusion models, such as Wan2.1, which are designed to support joint training on both video and image data. In particular, given an input video $V\in\mathbb{R}^{(1+T)\times H\times W\times 3}$, where the frames of $V$ follow the $1+T$ input format, Wan2.1 divides the video into $1+T/4$ chunks. Then, Wan-VAE compresses the spatio-temporal dimensions of these chunks to $[1+T/4,H/8,W/8]$, while the first frame is only spatially compressed to better handle image data. This independent processing of the first frame tends to cause forgetting and drifting issues.
@@ 3.2 Partial-update Training and DPO
After completing the first LoRA-based training stage, we obtain a diffusion transformer capable of seamless long video generation. Next, we introduce the Audio Cross-Attention module and adopt the Flow Matching (lipman2022flow) approach used in Wan2.1 to update its parameters. Specifically, we aggregate every four consecutive audio frames into a single representation to ensure temporal alignment between the audio features and the compressed video latent representation. The Audio Cross-Attention mechanism is defined as:
@@
To further improve lip-sync accuracy and align the model with human preferences, we introduce DPO for optimization after completing the aforementioned stages. Hallo4 (cui2025hallo4) presents the first audio-driven portrait DPO dataset that captures human preferences in lip-sync and facial naturalness via annotator rankings. Unlike Hallo4, which relies on human annotators to construct the dataset, we introduce DPO in a more efficient and cost-effective manner.
Direct Preference Optimization formulates the alignment of models with human preferences as a policy optimization task, based on pairwise preference data $\mathcal{D}=\{(x,y^{w},y^{l})\}$, where $y^{w}$ is preferred over $y^{l}$. As shown in Figure 2, for each training sample, we randomly select five segments and employ LatentSync (li2024latentsync) to compute the Sync-C score for each; the highest-scoring segment is selected as $y^{w}$, and the lowest as $y^{l}$. Finally, we use the Flow-DPO loss proposed by VideoReward (liu2025improving) to train the model:
@@ 3.3 Mask-CFG for Multi-Character Audio-driven Animation
After the training stage described above, we obtain a diffusion transformer that achieves accurate lip-sync and strong alignment with human preferences. We now introduce a training-free method to enable multi-character audio-driven video generation. Methods such as MultiTalk (kong2025let) and HunyuanVideo-Avatar achieve this capability by constructing multi-speaker datasets and modifying the cross-attention mechanism. In contrast, we enable multi-character animation by improving the classifier-free guidance (CFG) mechanism during inference—without any training or model modification—resulting in a simple, efficient, and framework-agnostic approach. Specifically, we propose Mask-CFG, which leverages spatial masks to route audio conditions to specific characters. Given an audio condition set $A=\{a_{1},a_{2},\dots,a_{n}\}$, the corresponding binary mask set is defined as $M=\{m_{1},m_{2},\dots,m_{n}\}$. Here, $a_{1}$ is considered to be silent audio, and $m_{1}$ serves as the background mask. Each $m_{i}\in\{0,1\}^{H\times W}$ denotes a binary segmentation of the input image, and the masks are exhaustive and mutually exclusive, satisfying $\bigvee_{i=1}^{n}m_{i}=\mathbf{1}$, meaning their union covers the entire image region. Under classifier-free guidance, the conditional distribution $p(a_{i}\mid x_{t})$ leads to the following formulation:
@@ 4.1 Experimental Setups
We collect our training data from public datasets (including AVSpeech (ephrat2018looking) and OpenHumanVid (li2025openhumanvid)) and sources we collect ourselves. To ensure high data quality, we employ tools such as Koala-36M (wang2025koala) to filter out videos with low brightness or poor aesthetic quality. Through this standardized selection process, we obtain over $300,000$ training samples, with a total duration exceeding $800$ hours. To demonstrate the effectiveness of our multi-character audio-driven approach in a training-free manner, all training samples are single-person talking videos. For evaluation, we use two public datasets: CelebV-HQ (zhu2022celebv), which features diverse scenes, and HDTF (zhang2021flow), which provides high-resolution videos and a larger number of subjects, to assess the animation capabilities of our method.
@@
The training weights of our first LoRA stage are initialized from the pretrained Wan2.1-I2V-14B-720P model, and the training process is conducted using $16$ NVIDIA A100 GPUs for a total of $5000$ steps. Subsequently, we use $32$ NVIDIA A100 GPUs and conduct additional training for $100,000$ steps to obtain the first $v_{\text{ref}}$. Next, we introduce DPO-based refinement training for another $100,000$ steps, during which $v_{\text{ref}}$ is updated every $10,000$ steps. The model operates at a resolution of $720\times 1280$. We employ AdamW as the optimizer and set the learning rate to $1\times 10^{-5}$. After completing the above training pipeline, we introduce Mask-CFG during the inference stage to enable multi-person audio-driven animation, with $\lambda$ set to $5.0$.